# Paseo de la Alameda 7, 46010, Valencia, Spain

### v1

@@ 4 Compton scattering phenomenology
$\frac{d\sigma_{e\gamma\to e\gamma}}{d\Omega}$ $=\frac{e^{4}\,Q^{4}}{64\,\pi^{2}s(1+\cos\theta)}\,\Big{(}2\cos\theta+\cos^{2}\theta+5+4\beta s(\cos^{2}\theta-1)\Big{)}$
@@
where $Q_{e}=-1$ is the charge of the electron and $e$ is the QED coupling constant. The full expression of the differential cross section including mass terms, is given in Appendix B.
@@
Figure 6: The deviation between the Standard Model and the GUP prediction $\Delta\sigma$ as a function of $\cos\theta$ for $\sqrt{s}=100$ (blue) and $200$ GeV (orange) for the two lower bounds on $\beta$ obtained in (37).
@@ 5 Conclusions
Future work will focus on extending the effective GUP-QED framework to other processes of phenomenological interest such as Higgs decays, electron and muon $(g-2)$ or loop-level corrections. Furthermore, incorporating renormalization group effects and developing a systematic approach to higher-order $\beta$ corrections could refine the predictions and improve the robustness of the bounds.
Appendix A Calculation of the Modified Feynman Propagators
@@
In order to find the Feynman propagator of the photon field, let us take a closer look at the photon kinetic terms from (3). If we chose a an appropriate gauge fixing term, i.e., [^5] One can easily check that the following expression, except the gauge-fixig term, is gauge-invariant.
$\mathcal{L}^{\text{GUP}}_{A,\xi}=-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+\frac{\beta}{2}F^{\mu\nu}\Big{(}\square F_{\mu\nu}\Big{)}-\frac{1}{2\xi}\Big{(}\partial_{\mu}(1-\beta\square)A^{\mu}\Big{)}^{2}\,,$ (44)
@@
$\left(g^{\mu\nu}\square(1-\beta\square)^{2}+\left(\frac{1}{\xi}-1\right)\partial^{\mu}\partial^{\nu}(1-\beta\square)^{2}\right)A_{\nu}\equiv O^{\mu\nu}_{\xi}A_{\nu}=0\,.$ (45)
@@
$\left(g^{\mu\nu}\left(\square(1-\beta\square)^{2}+m^{2}\right)-\partial^{\mu}\partial^{\nu}(1-\beta\square)^{2}\right)V_{\nu}\equiv O_{V}^{\mu\nu}V_{\nu}=0\,,$ (49)
@@
$\frac{1}{p^{2}-m^{2}}\left(-g_{\mu\nu}+\frac{p_{\mu}p_{\nu}}{m^{2}}\right)+\frac{2\beta\,p^{4}}{\left(p^{2}-m^{2}\right)^{2}}\left(g_{\mu\nu}-\frac{p_{\mu}p_{\nu}}{p^{2}}\right)+\mathcal{O}\left(\beta^{2}\right)\,.$ (51)
Appendix B Full Expression of the Compton Cross Section
The unpolarized, spin averaged, squared amplitude at $\mathcal{O}(\beta)$ for the process is given by the sum
|¯$|\overline{\mathcal{M}}|^{2}=\frac{1}{4}\left(\,|{\mathcal{M}}_{\text{SM}}|^{2}+2\,\text{Re}[\mathcal{M}_{\text{SM}}^{\dagger}\,\mathcal{M}_{\text{GUP}}]\,\right),$ (52)
@@ Appendix B Full Expression of the Compton Cross Section
$\qquad\qquad\qquad\qquad+m^{2}({s}+{u})\left({s}^{2}+6{s}{u}+{u}^{2}\right)-{s}{u}\left({s}^{2}+{u}^{2}\right)\Big{)}$ (53)
@@
$2\,\text{Re}[\mathcal{M}_{\text{SM}}^{\dagger}\,\mathcal{M}_{\text{GUP}}]$ $=\frac{16\beta e^{4}Q^{4}}{\left(m^{2}-{s}\right)^{3}\left(m^{2}-{u}\right)^{3}}\Big{(}-2{s}^{3}{u}^{3}({s}+{u})+22m^{14}-2m^{12}(7{s}+9{u})$
$\qquad+m^{10}\left(-5{s}^{2}-46{s}{u}+5{u}^{2}\right)+m^{8}\left(7{s}^{3}+51{s}^{2}{u}+35{s}{u}^{2}-{u}^{3}\right)$
$\qquad-2m^{6}{s}\left({s}^{3}+6{s}^{2}{u}-2{s}{u}^{2}-2{u}^{3}\right)+m^{4}{s}{u}\left({s}^{3}-23{s}^{2}{u}-31{s}{u}^{2}-3{u}^{3}\right)$
@@
Acknowledgements
We would like to thank A. Pich for helpful comments on this manuscript.

### v2

@@ 3 Compton scattering phenomenology
$\frac{d\sigma_{e\gamma\to e\gamma}}{d\Omega}$ $=\frac{e^{4}\,Q^{4}}{64\,\pi^{2}s(1+\cos\theta)}$
$\qquad\times\,\Big(2\cos\theta+\cos^{2}\theta+5+4\beta s(\cos^{2}\theta-1)\Big)$
@@
where $Q_{e}=-1$ is the charge of the electron, $e$ the QED coupling constant, and $\alpha$, the fine structure constant. The full expression of the differential cross section including mass terms, is given in Appendix C.
@@
((a))
((b))
Figure 6: Deviation between the Standard Model and the GUP prediction $\Delta\sigma$ as a function of $\cos\theta$ for two energy values, using the lower bounds on $\beta$ from Eq. (38).
@@ 4 Conclusions
Future work will focus on extending the effective GUP-QED framework to other processes of phenomenological interest such as Higgs decays, electron and muon (g2)(g-2) or loop-level corrections. Furthermore, incorporating renormalization group effects and developing a systematic approach to higher-order $\beta$ corrections could refine the predictions and improve the robustness of the bounds.
Appendix A Deformed phase space, Poincaré generators and scattering framework
In this study we assume a commutative space-time and commuting physical momenta, [^4] We will eliminate the hat notation for quantum operators for simplicity.
$[x^{\mu},x^{\nu}]$ $=0,$ (40)
$[p^{\mu},p^{\nu}]$ $=0.$ (41)
The deformation is introduced through the covariant generalized commutation relation
$[x^{\mu},p^{\nu}]$ $=-i\eta^{\mu\nu}(1+\beta p^{2})-2i\beta p^{\mu}p^{\nu},$ (42)
where
$p^{2}\equiv\eta_{\mu\nu}p^{\mu}p^{\nu},$ (43)
and $\beta=\beta_{0}\,\Lambda^{-2}$ is the deformation parameter. The standard Heisenberg algebra is recovered in the limit $\beta\to 0$. Space-time translations are generated by the physical momentum operator,
$T(a)=\exp(-ia_{\mu}p^{\mu}).$ (44)
Using the Baker–Campbell–Hausdorff expansion, the action of a translation on an operator $A$ is
$T(a)\,A\,T^{\dagger}(a)$ $=A+i[-a_{\rho}p^{\rho},A]$
$\qquad+\frac{i^{2}}{2!}[-a_{\rho}p^{\rho},[-a_{\sigma}p^{\sigma},A]]+\cdots.$ (45)
Applying this to the spacetime coordinates and using the deformed algebra (42), one finds
$i[-a_{\rho}p^{\rho},x^{\mu}]$ $=a^{\mu}(1+\beta p^{2})+2\beta(a_{\rho}p^{\rho})\,p^{\mu}.$ (46)
Since the physical momenta commute, $p^{\mu},p^{\nu}]=0$, it follows that
$[-a_{\rho}p^{\rho},\,(1+\beta p^{2})]=0,\qquad[-a_{\rho}p^{\rho},\,(a_{\rho}p^{\rho})p^{\mu}]=0,$ (47)
so that all higher nested commutators vanish and the translation action closes at first order,
$T(a)\,x^{\mu}\,T^{\dagger}(a)$ $=x^{\mu}+a^{\mu}(1+\beta p^{2})+2\beta(a_{\rho}p^{\rho})\,p^{\mu},$ (48)
$T(a)\,p^{\mu}\,T^{\dagger}(a)$ $=p^{\mu}.$ (49)
On the other hand, for a multiparticle system, the physical momentum acts additively on the tensor-product Hilbert space. For instance, in a two-particle system,
$p^{\mu}_{(1)}$ $\equiv p^{\mu}_{1}\otimes\mathbf{1},\qquad p^{\mu}_{(2)}\equiv\mathbf{1}\otimes p^{\mu}_{2},$ (50)
with the property
$[p^{\mu}_{(1)},p^{\nu}_{(2)}]=0,$ (51)
we obtain that the total translation operator factorizes, i.e.,
$T_{\rm tot}(a)$ $)\displaystyle=\exp\!\big(-ia_{\mu}(p^{\mu}_{(1)}+p^{\mu}_{(2)}$$\big)$
[…]
and the generator of translations is the total momentum
$p^{\mu}_{\rm tot}=\sum_{i}p^{\mu}_{(i)}.$ (53)
Translation invariance of the $S$-matrix therefore implies standard momentum conservation in scattering processes,
$\sum_{\rm in}p^{\mu}=\sum_{\rm out}p^{\mu}.$ (54)
Let us now, finally, consider the Lorentz generators of the modified algebra. The standard expression for $M^{\mu\nu}$ given by
$M^{\mu\nu}_{0}=x^{\mu}p^{\nu}-x^{\nu}p^{\mu}$ (55)
does not generate standard Lorentz transformations on $p^{\mu}$ when the deformed commutator (42) is used. To restore the standard Lorentz action on physical momentum, we define the modified generators
$\widetilde{M}^{\mu\nu}$ $\equiv\frac{1}{1+\beta p^{2}}\,\big(x^{\mu}p^{\nu}-x^{\nu}p^{\mu}\big)$
$=M^{\mu\nu}_{0}(1-\beta p^{2})+\mathcal{O}(\beta^{2}).$ (56)
Using (42), one finds, at $\mathcal{O}(\beta)$
$[\widetilde{M}^{\mu\nu},p^{\rho}]$ $=i\big(\eta^{\nu\rho}p^{\mu}-\eta^{\mu\rho}p^{\nu}\big),$ (57)
so that $p^{\mu}$ transforms as a standard Lorentz four-vector. In particular,
$[\widetilde{M}^{\mu\nu},p^{2}]=0,$ (58)
therefore, if we impose the mass shell-condition
$p^{2}=m^{2},$ (59)
we can observe that it is a Casimir and Lorentz invariant quantity. The action of $\widetilde{M}^{\mu\nu}$ on $x^{\rho}$ is momentum dependent,
$[\widetilde{M}^{\mu\nu},x^{\rho}]$ $=i(\eta^{\nu\rho}x^{\mu}-\eta^{\mu\rho}x^{\nu})+\mathcal{O}(\beta),$ (60)
so spacetime coordinates do not transform as a pure Lorentz four-vector beyond the $\beta\to 0$ limit. This reflects the fact that Lorentz symmetry is realized nonlinearly in position space, while remaining exact in momentum space. Such behavior should not come as a surprise in models inspired by the existence of a minimal length, where the operational meaning of spacetime coordinates is known to be modified and localization properties become momentum dependent. In this context, $x^{\mu}$ should be regarded as an auxiliary phase-space variable, whereas physical observables and scattering processes are most naturally formulated in momentum space.
Since $p^{\mu}$ transforms as a four-vector and $p^{2}$ is a Lorentz scalar, the right-hand side of (42) is manifestly covariant. A finite Lorentz transformation is implemented by the unitary operator
$U(\Lambda)$ $=\exp\!\left(-\frac{i}{2}\,\omega_{\mu\nu}\,\widetilde{M}^{\mu\nu}\right),$ (61)
where $\omega_{\mu\nu}=-\omega_{\nu\mu}$ are the real (infinitesimal) Lorentz parameters. To first order, the Lorentz matrix is
$\Lambda^{\mu}{}_{\nu}$ $=\delta^{\mu}{}_{\nu}+\omega^{\mu}{}_{\nu}+\mathcal{O}(\omega^{2}),\qquad\omega^{\mu}{}_{\nu}\equiv\eta^{\mu\rho}\omega_{\rho\nu},$ (62)
and Lorentz invariance implies
$\Lambda^{T}\eta\Lambda=\eta\quad\Longleftrightarrow\quad\omega_{\mu\nu}=-\omega_{\nu\mu}.$ (63)
Under the previous transformations, the transformed operators are given by
$x^{\prime\mu}$ $=U(\Lambda)\,x^{\mu}\,U^{\dagger}(\Lambda),$ (64)
$p^{\prime\mu}$ $=U(\Lambda)\,p^{\mu}\,U^{\dagger}(\Lambda).$ (65)
Using the usual expansion,
$UAU^{\dagger}$ $=A+i\left[\frac{1}{2}\omega_{\rho\sigma}\widetilde{M}^{\rho\sigma},A\right]$
$\;\;+\frac{i^{2}}{2!}\left[\frac{1}{2}\omega_{\rho\sigma}\widetilde{M}^{\rho\sigma},\left[\frac{1}{2}\omega_{\alpha\beta}\widetilde{M}^{\alpha\beta},A\right]\right]+\cdots,$ (66)
one obtains the infinitesimal variations
$\delta x^{\mu}\equiv x^{\prime\mu}-x^{\mu}$ $=i\left[\frac{1}{2}\omega_{\rho\sigma}\widetilde{M}^{\rho\sigma},x^{\mu}\right]+\mathcal{O}(\omega^{2}),$ (67)
$\delta p^{\mu}\equiv p^{\prime\mu}-p^{\mu}$ $=i\left[\frac{1}{2}\omega_{\rho\sigma}\widetilde{M}^{\rho\sigma},p^{\mu}\right]+\mathcal{O}(\omega^{2}).$ (68)
By construction of $\widetilde{M}^{\mu\nu}$ (56), the momentum transforms as a standard Lorentz four-vector,
$[\widetilde{M}^{\rho\sigma},p^{\mu}]$ $=i\left(\eta^{\sigma\mu}p^{\rho}-\eta^{\rho\mu}p^{\sigma}\right),$ (69)
and therefore
$\delta p^{\mu}$ $=\omega^{\mu}{}_{\nu}\,p^{\nu}+\mathcal{O}(\omega^{2}),\qquad p^{\prime\mu}=\Lambda^{\mu}{}_{\nu}p^{\nu}+\mathcal{O}(\omega^{2}).$ (70)
Finally, the transformed commutator can be written explicitly in terms of the generator as
$[x^{\prime\mu},p^{\prime\nu}]$ $=\big[Ux^{\mu}U^{\dagger},\;Up^{\nu}U^{\dagger}\big]=U\,[x^{\mu},p^{\nu}]\,U^{\dagger}$ (71)
$=[x^{\mu},p^{\nu}]+i\left[\frac{1}{2}\omega_{\rho\sigma}\widetilde{M}^{\rho\sigma},[x^{\mu},p^{\nu}]\right]+\mathcal{O}(\omega^{2}).$
Using the deformed algebra (42), this implies the form-invariant result
$[x^{\prime\mu},p^{\prime\nu}]$ $=-i\Big(\eta^{\mu\nu}(1+\beta p^{\prime 2})+2\beta p^{\prime\mu}p^{\prime\nu}\Big),$ (72)
with $p^{\prime 2}=\eta_{\rho\sigma}p^{\prime\rho}p^{\prime\sigma}=p^{2}$.
Because translations are generated by the additive physical momentum $p^{\mu}$, scattering amplitudes satisfy standard momentum conservation,
$(2\pi)^{4}\delta^{(4)}\!\left(\sum_{\rm in}p^{\mu}-\sum_{\rm out}p^{\mu}\right).$ (73)
External states are placed on the physical mass shell (59), and Lorentz invariance of amplitudes is ensured by building interaction vertices from Lorentz scalars constructed out of $p^{\mu}$. All deformation effects can therefore be encoded in higher-dimensional, momentum-dependent operators suppressed by $\Lambda$, while standard momentum-space scattering theory remains applicable within the regime
$|p^{2}|\ll\Lambda^{2}.$ (74)
Appendix B Calculation of the Modified Feynman Propagators
@@
[…]
In order to find the Feynman propagator of the photon field, let us take a closer look at the photon kinetic terms from (2). If we chose a an appropriate gauge fixing term, i.e., (one can easily check that the following expression, except the gauge-fixig term, is gauge-invariant)
$\mathcal{L}^{\text{GUP}}_{A,\xi}$ $=-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+\frac{\beta}{2}F^{\mu\nu}\Big(\square F_{\mu\nu}\Big)$
$\qquad-\frac{1}{2\xi}\Big(\partial_{\mu}(1-\beta\square)A^{\mu}\Big)^{2}\,$, (80)
@@
$O^{\mu\nu}_{\xi}A_{\nu}=0\,.$ (81)
with $O^{\mu\nu}_{\xi}$ defined as
$O^{\mu\nu}_{\xi}\equiv\left(g^{\mu\nu}\square(1-\beta\square)^{2}+\left(\frac{1}{\xi}-1\right)\partial^{\mu}\partial^{\nu}(1-\beta\square)^{2}\right).$ (82)
@@
$O_{V}^{\mu\nu}V_{\nu}=0,$ (86)
with $O_{V}^{\mu\nu}$
$O_{V}^{\mu\nu}\equiv\Big(g^{\mu\nu}\left(\square(1-\beta\square)^{2}+m^{2}\right)-\partial^{\mu}\partial^{\nu}(1-\beta\square)^{2}\Big)\,$, (87)
@@
$\frac{1}{p^{2}-m^{2}}\left(-g_{\mu\nu}+\frac{p_{\mu}p_{\nu}}{m^{2}}\right)+$
$\hskip 18.49988pt\frac{2\beta\,p^{4}}{\left(p^{2}-m^{2}\right)^{2}}\left(g_{\mu\nu}-\frac{p_{\mu}p_{\nu}}{p^{2}}\right)+\mathcal{O}\left(\beta^{2}\right)\,.$ (89)
Appendix C Full Expression of the Compton Cross Section
The unpolarized, spin-averaged, squared amplitude at $\mathcal{O}(\beta)$ for the process is given by the sum
$|\overline{\mathcal{M}}|^{2}=\frac{1}{4}\left(\,|{\mathcal{M}}_{\text{SM}}|^{2}+2\,\text{Re}[\mathcal{M}_{\text{SM}}^{\dagger}\,\mathcal{M}_{\text{GUP}}]\,\right),$ (90)
@@ Appendix C Full Expression of the Compton Cross Section
$\;\;+m^{2}({s}+{u})\left({s}^{2}+6{s}{u}+{u}^{2}\right)-{s}{u}\left({s}^{2}+{u}^{2}\right)\Big)$ (91)
@@
$2\,\text{Re}\left[\mathcal{M}_{\text{SM}}^{\dagger}\,\mathcal{M}_{\text{GUP}}\right]$ $=\frac{16\beta e^{4}Q^{4}}{(m^{2}-s)^{3}(m^{2}-u)^{3}}\Big(-2s^{3}u^{3}(s+u)$
$\qquad+22m^{14}-2m^{12}(7s+9u)$
$\qquad+m^{10}\left(-5s^{2}-46su+5u^{2}\right)$
$\qquad+m^{8}\left(7s^{3}+51s^{2}u+35su^{2}-u^{3}\right)$
$\qquad-2m^{6}s\left(s^{3}+6s^{2}u-2su^{2}-2u^{3}\right))$
$\qquad+m^{4}su\left(s^{3}-23s^{2}u-31su^{2}-3u^{3}\right)$